# TDDBench: A Benchmark for Training data detection

### v1

@@ A.1.1 Metric-based dectection
A.1.2 Learning-based dectection
@@
A.1.3 Model-based dectection
@@
A.1.4 Query-based dectection
@@
A.2 Training data detection algorithms in large language models
In this section, we provide a brief introduction to the TDD algorithms for large language models; please refer to the related works for specific details.
MIN-K(Shi et al., 2024) utilizes the kof tokens with the lowest likelihoods as the detection basis, rather than average loss.
MIN-K++ (Zhang et al., 2024) is based on the observation that training samples tend to be local maxima of the modeled distribution along each input dimension due to maximum likelihood training, effectively translating the TDD problem into the identification of local maxima.
A.3 Performances and training details of target models
@@
A.4 Detection performance with more queries and reference models.
In this part, we analyze the potential of the two latest types of TDD algorithms: reference-based and query-based. Specifically, we expand the number of reference models that the reference-based algorithm can create and the number of queries that the query-based algorithm can handle. Due to limitations in computational resources, we opt to decrease rather than increase the number of reference models in the text modality. We anticipate that similar conclusions can be drawn with a larger number of reference models. The results in Figure 4 demonstrate that increasing the number of queries or reference models can improve the TDD algorithm’s performance. Adjusting the number of reference models has a more significant impact on the algorithm’s performance compared to increasing the number of queries. Nevertheless, enhancing computational resources does not close the gap in algorithm design. For instance, in tabular data, even with 256 reference models, other algorithms fail to outperform LiRA with 16 reference models.
(a) Image
(b) Tabular
(c) Text
(d) Image
(e) Tabular
(f) Text
Figure 4: AUROC of TDD algorithms with more queries and reference models. The abbreviations Query-nei, Q-ref, M-loss, M-cal, M-lira, M-fpr, and M-robust stand for Query-neighbor, Query-ref, Model-loss, Model-calibration, Model-lira, Model-fpr, and Model-robust, respectively. While an increase in available computing resources can enhance TDD performance to some extent, the improvement is not significant. Therefore, a more cost-effective approach is to focus on designing more powerful algorithms.
A.5 Complete version of the experimental results
@@
A.6 Performance under different metrics

### v2

@@ A.6.1 Metric-based dectection
A.6.2 Learning-based dectection
@@
A.6.3 Model-based dectection
@@
A.6.4 Query-based dectection
@@
A.6.5 How to operate a Model-based or Query-based TDD algorithms
In this section, we outline the implementation of Model-based and Query-based algorithms. Specifically, we demonstrate how to train a reference model for focal data $x$ in the Model-based algorithms, as well as how to obtain additional query results in the Query-based algorithms. By following these steps in Alg 1 and Alg 2, you can effectively implement both Model-based and Query-based TDD algorithms.
Input: Reference dataset $D$, focal data $x$, target model $f$, number of reference models NN;
Output: Whether $x$ was used to train $f$
1 for NN times do
2   Sample a subset from $\mathbb{D}$ ;
3   Train a reference model using the combined dataset $\mathbb{D}\cup d$ ;
4   Obtain $x$’s detection metric (e.g. loss) from this reference model, which is trained with $x$ ;
5   Train a reference model using the dataset $\mathbb{D}\setminus d$ ;
6   Obtain $x$’s detection metric (e.g. loss) from this reference model, which is trained without $x$
7 end for
8Obtain $x$’s detection metric from the target model $f$ ;
9 Implement Model-based TDD using the detection criterion from reference models and target model ;
Algorithm 1 How to train reference models in Model-based TDD algorithms
Input: Focal data $x$, target model $f*$, number of queries per sample NN;
Output: Whether $x$ was used to train $f$
1 for NN times do
2   Modify the data point $x$ based on the chosen data augmentation strategy (e.g., add noise, flip) ;
3   Input the modified data $d^{\prime}$ into the target model $f*$ to obtain the query results
4 end for
5Implement training data detection using the query results obtained from the different modified data points ;
Algorithm 2 How to obtain extra queries in Query-based TDD algorithms
A.7 Performances and training details of target models
In Table 13, we present the accuracy of the target model across various datasets. Table 14 extends this evaluation by showcasing the accuracy of the target model under diverse model architectures. Additionally, Table 15 complements these results by detailing the training configurations and hyperparameters used for each model, offering transparency and reproducibility for the experimental setup.
@@ A.7 Performances and training details of target models
A.8 Complete version of the experimental results
The comprehensive evaluation of TDD performance across different datasets, various target model architectures, and diverse shadow and reference models is presented in Table 16, Table 17, and Table 18, respectively.
@@ A.8 Complete version of the experimental results
A.9 Performance under different metrics
We present the TDD performance evaluated across multiple metrics for three distinct data modalities, as detailed in Table 19, Table 20, and Table 21.